OpenStudy (anonymous):

(medal for the person who helps) Use the slope formula to find the missing coordinate. R(0, _), S(-4, 2) m = 5 A.)18 B.)22 C.)-22

OpenStudy (anonymous):

B. 100% sure

OpenStudy (anonymous):

hey buddy, you ready to learn? xD

OpenStudy (anonymous):

can you guys explain how to get the answer you think it is ? cuz I don't under stand slope at all

OpenStudy (anonymous):

thats exactly what i was gonna do

OpenStudy (anonymous):

thanks ~

OpenStudy (anonymous):

so how do i do it ?

OpenStudy (anonymous):

okay so we can do: m=y2-y1/x2-x1 so 5=2-y/-4-0 solve for y :)

OpenStudy (anonymous):

ok, so can you draw out how to solve for y ?

OpenStudy (anonymous):

yea give me a sec :)

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

so multiply both sides by the denominator: